OpenStudy (anonymous):

Help me find Q please!! For mutually exclusive events R1, R2, and R3, we have P(R1)=0.05, P(R2)=0.4, P(R3)=0.55. Also P(Q|R1)=0.5, P(Q|R2)=0.7, and P(Q|R3)=0.4. Find P(R2|Q).

OpenStudy (xapproachesinfinity):

P(R2/Q)=P(R2 and Q)/P(Q) we have P(Q/R2)=P(R2 and Q)/P(R2) hence P(R2 and Q)= 0.7P(R2)=0.7x0.4=0.28

OpenStudy (xapproachesinfinity):

we need P(Q) now!

OpenStudy (anonymous):

isnt 1-P=Q?

OpenStudy (xapproachesinfinity):

P of what? p here means probability! so probability of what event are you talking about

OpenStudy (anonymous):

1-P(R2)?

OpenStudy (xapproachesinfinity):

I think there a formula that link P(A/B) with P(B/A) sort of thing but i forget

OpenStudy (xapproachesinfinity):

that will be true if Q is complement of R2 but that's not true?

OpenStudy (anonymous):

ahh idk im just guessing, i'm not too good at this :/

OpenStudy (xapproachesinfinity):

We need that formula to get this done! and used the fact the those events are mutually exclusive!

OpenStudy (xapproachesinfinity):

use*

OpenStudy (anonymous):

ill look for it!

OpenStudy (xapproachesinfinity):

@SolomonZelman probab hehe

OpenStudy (xapproachesinfinity):

@nincompoop